Dad is seven times as old as his son. 10 years later, he will be three times as old as his son. Find their ages.
Answer by jim_thompson5910(35256)

"Dad is seven times as old as his son" ---> D = 7S

"10 years later, he will be three times as old as his son" ---> D+10 = 3(S+10)


D+10 = 3(S+10)

7S+10 = 3(S+10)

7S+10 = 3S+30

7S = 3S+30-10

7S-3S = 30-10

4S = 20

S = 20/4

S = 5 <--- The son is 5 years old

D = 7S

D = 7*5

D = 35 <---- The dad is 35 years old.